OpenStudy (anonymous):

Write an exponential function whose graph passes (0,1) and (6,64)?

OpenStudy (anonymous):

you need to find: \[e^{f(x)} \] such that\[e^{f(0)} = 1\] and \[e^{f(6)} = 64\] You can pick any function that satisfies those. Simplest function to take would be a linear function, and in this case we can: f(x) = ax + b and we want to find a and b that satisfy these; and we don't even have to use b, but we can. (they said pick one function. i can come up with an infinite amount) 6a = 64 --> a = 32/3 so a function that works is: \[e^{\frac{ 32x }{ 3 }}\]

OpenStudy (anonymous):

Thank U

OpenStudy (anonymous):

but I dont understand why u say linear

OpenStudy (anonymous):

because the question ask for exponential function

OpenStudy (anonymous):

I need clarification as a blueprint to do the many others that I have in this regard... HELP

OpenStudy (anonymous):

by linear i meant the f(x). in e^(f(x)) for f(x) I could have picked a non-linear function (2nd order or higher), but I would have just complicated things. I could have found a function ax^2 + bx that would have satisfied the two conditions. But then again that's complicating things. You can always solve this using a linear function (a first order polynomial), which looks like ax + b (where b can = 0) exists in the form: \[e^{ax + b} \] or simply and often: \[e^{ax}\] I hope i cleaned up my mess a bit :P

OpenStudy (anonymous):

wow my answer is wrong above. very wrong

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

e^(ax)=64=e^(6a) so to solve for a the real actual way (idk what i was thinking) ln(64) = 6a a = ln64/6 function is then: \[f(x) = e^{\frac{ (\ln64)x} {6 }}\]

OpenStudy (anonymous):

OK thank U